OpenStudy (anonymous):

How do you simplify this? n^2-4/n+2 multiplied by n/3n-6

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (angela210793):

n^2-4 can be written as =(n-2)*(n+2) so we'll have (n-2)*(n+2) ------------= n-2 (n+2)

OpenStudy (anonymous):

I need to simplify the whole equation. So I would have to multipy that to n/3n-6

OpenStudy (angela210793):

oops sorry i 4got the rest of the problem

OpenStudy (anonymous):

It's okay.

OpenStudy (angela210793):

(n-2)*n/3n-6=(n^2-2n)/(3n-6)

OpenStudy (anonymous):

Is that the whole answer? I'm looking for just one answer that is simplified. I don't know the name of this kind of math method though...

OpenStudy (anonymous):

Maybe an example would help?

OpenStudy (dumbcow):

3n -6 = 3(n-2) \[\frac{n(n-2)}{3(n-2)} = \frac{n}{3}\]

OpenStudy (anonymous):

For example, y^2/12x is in Simplest Form

OpenStudy (dumbcow):

the math method is just factoring and reducing by cancelling

OpenStudy (anonymous):

Yes, that's it!

OpenStudy (anonymous):

I don't understand how to do the problem.

OpenStudy (dumbcow):

break it down, look at each term and try to factor it then cancel identical terms on top and bottom

OpenStudy (dumbcow):

best way to learn is to practice factoring

OpenStudy (anonymous):

OKay, I saw your above answer, thank you! I think i can figure it out myself now